# Computing Lower and Upper Hitting Probabilities for Imprecise Markov Chains

### v1

@@ 3.1 Lower and Upper Reachability
For all $T\in\mathcal{T}$ there exists $n\in\mathbb{N}$ (either 1 or 2) such that $T^{n}(1,3)>0$. Hence, $3$ is lower reachable from $1$, under $\operatorname{(LR0)}$. However, it is not true that there exists $n\in\mathbb{N}$ such that $T^{n}(1,3)>0$ for all $T\in\mathcal{T}$. In fact, the matrix $T_{1}\in\mathcal{T}$, defined by fixing $T_{1}(1,2)=p=1$, has $T_{1}^{2}(1,3)=1>0$ and $T_{1}^{1}(1,3)=0$. On the other hand, the matrix $T_{0}\in\mathcal{T}$, defined by fixing $T_{0}(1,2)=p=0$, has $T_{0}^{1}(1,3)=1$ and $T_{0}^{2}(1,3)=0$. So $3$ is not lower reachable from $1$, under $\operatorname{(LR2)}$, because for all $n\in\mathbb{N}_{0}$ there exists $T\in\mathcal{T}$ such that $T^{n}\mathbb{1}_{\{3\}}](1)=0$.
@@
If $n=6$, all $T\in\mathcal{T}$ are such that $T^{6}(1,3)>0$, i.e. in six steps, whichever $T$ is selected, the process has a positive probability of being in state 3, if it has started in state 1. Therefore, 3 is lower reachable from 1, under $\operatorname{(LR2)}$. However, for all $n\in\mathbb{N}$ we have
@@
implying that 3 is not lower reachable from 1, under $\operatorname{(LR3)}$. This comes from the fact that, since we are applying the lower transition operator $\underline{T}$ $n$ times, there is no guarantee that the same transition matrix achieves the minimum. Given $n\in\mathbb{N}$, we can find a sequence of $n$ transition matrices that minimises the probability of hitting state 3 in exactly $n$ steps, and we find that this probability is zero. Given $n\in\mathbb{N}$, the sequence of transition matrices $(T_{k})_{k\in\mathbb{N}}$ defined as
$T_{k}\coloneqq\begin{cases}S_{0},&\text{ if $n$ is odd and $k=1$},\\ S_{1},&\text{otherwise},\end{cases}$
where $S_{0,1}\in\mathcal{T}$ are the two matrix defined by setting either $q=0$ or $q=1$, ensures to avoid state 3 at time $n$. Thus, 3 is not lower reachable from 1, under $\operatorname{(LR3)}$.
There are particular cases in which (LR0), (LR1), (LR2) and (LR3) are all equivalent. We say that the set $\mathcal{D}$ is closed when for all $x\in\mathcal{D},y\notin\mathcal{D}$ and for all $T\in\mathcal{T}$, we have $T(x,y)=0$.
@@
The fact that (LR0) $\Rightarrow$ (LR2) comes from the fact that (LR0) guarantees for all $T\in\mathcal{T}$ the existence of a time $n=n(T)$ such that $T^{n}\mathbb{1}_{\mathcal{D}}](x)>0$. Then $n^{*}\coloneqq\max_{T\in\mathcal{T}}n(T)$ is the time for which (LR2) is satisfied.
To prove that $x$ lower reaches $y$ under (LR3) if it does under (LR1), let us recall that we need to prove that there exists $n\in\mathbb{N}$ such that
@@
for all $T_{1},\dots,T_{n}\in\mathcal{T}$. Let us pick $n=k^{*}$ recalling that, by hypothesis, we have $x\in\mathcal{D}_{k^{*}}$. For all $T_{1}\in\mathcal{T}$, there exists $x_{1}\in\mathcal{D}_{k^{*}-1}$ such that $T_{1}(x,x_{1})>0$. Iterating this argument will get to $\mathcal{D}$ in at most $k^{*}$ steps. If, for some sequence of matrices, there is a positive probability of hitting $\mathcal{D}$ at time $k^{\prime}<k^{*}$, then, due to the closedness of $\mathcal{D}$, the same would hold at time $k^{*}$. ∎
@@ 3.2 Lower Hitting Probabilities
This set can be rewritten as
$=\{x\in A^{c}\mid\exists T\in\mathcal{T}:x\in\mathcal{C}_{T}\}.$
We may interpret $\mathcal{A}_{\mathcal{T}}$ as the set of states whose lower probability of reaching $A$ is zero, since there exists a transition matrix that disconnects $\mathcal{A}_{\mathcal{T}}$ from $A$. Using the definition of lower reachability, we deduce a straightforward relationship between the sets $\mathcal{A}_{\mathcal{T}}$ and $\mathcal{C}_{T}$.
@@
The proof follows immediately from the definition of $\mathcal{A}_{\mathcal{T}}$ and lower reachability. ∎
We are finally able to characterise lower hitting probabilities as the unique solution to a system of equations.
@@
The vector of lower hitting probabilities $\underline{p}^{\mathcal{T}}$ satisfies
1.
$\underline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$,
2.
$\underline{p}^{\mathcal{T}}(x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$,
3.
$0<\underline{p}^{\mathcal{T}}(x)\leq 1$ for all $x\in A^{c}\setminus\mathcal{A}_{\mathcal{T}}$,
and it is the only solution to
$\underline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\underline{T}\,\underline{p}^{\mathcal{T}}.$ (13)
satisfying conditions 1 and 2.
@@
This result is derived similarly to Krak’s Proposition 2 [krak2021comphit]. Let us start with the claims about the structure of $\underline{p}^{\mathcal{T}}$. We know from Krak et al. [krak2019hitting, Corollary 19] that Equation (13) has a minimal nonnegative solution $\underline{p}^{\mathcal{T}}\in[0,1]^{\mathcal{X}}$. By the definition of the system, this solution satisfies $\underline{p}^{\mathcal{T}}(x)=1$ for $x\in A$. The fact that $\underline{p}^{\mathcal{T}}(x)=0$ if and only if $x\in\mathcal{A}_{\mathcal{T}}$ follows from the definition of lower hitting probability, Lemma 3.8, and Proposition 2.2.
Now we prove the uniqueness of $\underline{p}^{\mathcal{T}}$. We want to show that if $p_{1},p_{2}\in[0,1]^{\mathcal{X}}$ are solutions of the system in Equation (13) and have $p_{1}(x)=p_{2}(x)=1$ for all $x\in A$ and $p_{1}(x)=p_{2}(x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$, then it must be $p_{1}=p_{2}$. Note that it suffices to show that $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$.
By Propositions 2.18, there exist $T_{1},T_{2}\in\mathcal{T}$ such that
$T_{1}p_{1}=\underline{T}p_{1}\quad\text{and}\quad T_{2}p_{2}=\underline{T}p_{2}.$
Since $p_{1}$ and $p_{2}$ are solutions to the system, it holds that
$p_{1}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{1}p_{1},$ (14)
and
$p_{2}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{2}p_{2}.$ (15)
Since $p_{1}$ and $p_{2}$ satisfy the systems in Equations (14) and (15), respectively, and both satisfy (13), we find that
$(T_{1}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=(\underline{T}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}\quad\text{ and }\quad(T_{2}p_{2})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=(\underline{T}p_{2})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}.$
We then get
$(T_{1}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=(\underline{T}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}\leq(T_{2}p_{1})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}.$
Equation (14) then yields
$p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$ $=(T_{1}p_{1})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$
$\leq(T_{2}p_{1})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$
$=T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\,p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}+(T_{2}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}},$
@@
where we applied Lemma 2.6 in the last step. Recursively expanding this expression for $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$ yields, after $m\in\mathbb{N}$ expansions:
$p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq(T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{m}p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}+\sum_{k=0}^{m-1}(T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{k}(T_{2}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}.$
Observe that $\mathcal{A}_{\mathcal{T}}=\bigcup_{T\in\mathcal{T}}\mathcal{C}_{T}$ by Lemma 3.8 and $\mathcal{C}_{T_{2}}\subseteq\mathcal{A}_{\mathcal{T}}\subset A^{c}$. Taking the limits as $m\to+\infty$ and applying Lemma 2.12 yields
$p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq(I-T_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}})^{-1}(T_{2}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}},$
where the last equality follows from Proposition 2.13. Therefore, without additional assumptions on $p_{1}$ and $p_{2}$, we have $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$. Similarly, by symmetry, $p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$, and thus $p_{1}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=p_{2}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$. ∎
@@ 3.3 Upper Hitting Probabilities
This set can be rewritten as
$=\{x\in A^{c}\mid\forall T\in\mathcal{T}:x\in\mathcal{C}_{T}\}.$
We may interpret $\mathscr{W}_{\mathcal{T}}$ as the set of states whose upper probability of reaching $A$ is zero, since every transition matrix disconnects $\mathscr{W}_{\mathcal{T}}$ to $A$. Similarly to lower hitting probabilities, there is a straightforward connection between sets $\mathscr{W}_{\mathcal{T}}$ and $\mathcal{C}_{T}$
@@
The proof follows immediately from the definition of $\mathscr{W}_{\mathcal{T}}$ and upper reachability. ∎
The following result characterises upper hitting probabilities as the unique solution to a system of equations.
@@
The vector of upper hitting probabilities $\overline{p}^{\mathcal{T}}$ satisfies
I.
$\overline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$,
II.
$\overline{p}^{\mathcal{T}}(x)=0$ for all $x\in\mathscr{W}_{\mathcal{T}}$,
III.
$0<\overline{p}^{\mathcal{T}}(x)\leq 1$ for all $x\in A^{c}\setminus\mathscr{W}_{\mathcal{T}}$,
and it is the only solution to
$\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\overline{T}\,\overline{p}^{\mathcal{T}}$ (17)
satisfying conditions I and II.

### v2

@@ 3.1 Lower and Upper Reachability
For all $T\in\mathcal{T}$ there exists $n\in\mathbb{N}$ (either 1 or 2) such that $T^{n}(1,3)>0$. Hence, $1$ lower reaches $3$, under $\operatorname{(LR0)}$. However, it is not true that there exists $n\in\mathbb{N}$ such that $T^{n}(1,3)>0$ for all $T\in\mathcal{T}$. In fact, the matrix $T_{1}\in\mathcal{T}$, defined by fixing $T_{1}(1,2)=p=1$, has $T_{1}^{1}(1,3)=0$, $T_{1}^{2}(1,3)=1$ and $T^{n}_{1}(1,3)=0$ for all $n\geq 3$. On the other hand, the matrix $T_{0}\in\mathcal{T}$, defined by fixing $T_{0}(1,2)=p=0$, has $T_{0}^{1}(1,3)=1$ and $T_{0}^{n}(1,3)=0$ for all $n\geq 2$. So $1$ does not lower reach $3$, under $\operatorname{(LR2)}$, because for all $n\in\mathbb{N}_{0}$ there exists $T\in\mathcal{T}$ such that $T^{n}\mathbb{1}_{\{3\}}](1)=0$.
@@
If $n=5$, all $T\in\mathcal{T}$ are such that $T^{5}(1,3)>0$, i.e. in five steps, whichever $T$ is selected, the process has a positive probability of being in state 3, if it has started in state 1. Therefore, 1 lower reaches 3, under $\operatorname{(LR2)}$. However, for all $n\in\mathbb{N}$ we have
@@
implying that 1 does not lower reach 3, under $\operatorname{(LR3)}$. This comes from the fact that, since we are applying the lower transition operator $\underline{T}$ $n$ times, there is no guarantee that the same transition matrix achieves the minimum. Given $n\in\mathbb{N}$, we can find a sequence of $n$ transition matrices that minimises the probability of hitting state 3 in exactly $n$ steps, and we find that this probability is zero. Given $n\in\mathbb{N}$, the sequence of transition matrices $(T_{k})_{k=1}^{n}$ defined as
$T_{k}\coloneqq\begin{cases}S_{0},&\text{if $k=n$},\\ S_{1},&\text{otherwise},\end{cases}$
where $S_{0,1}\in\mathcal{T}$ are the two matrix defined by setting either $q=0$ or $q=1$, ensures to avoid state 3 at time $n$. Thus, 1 does not lower reach 3, under $\operatorname{(LR3)}$. To check this, one may also observe that $\underline{T}\mathbb{1}_{\{3\}}=\mathbb{1}_{\{2\}}$ (the matrix $S_{0}$ achieves this minimum), $\underline{T}^{2}\mathbb{1}_{\{3\}}=\underline{T}\mathbb{1}_{\{2\}}=\mathbf{0}$ (the sequence $(S_{1},S_{0})$ achieves this minimum), thus
$\underline{T}^{n+2}\mathbb{1}_{\{3\}}](1)=[\underline{T}^{n+1}\mathbb{1}_{\{2\}}](1)=[\underline{T}^{n}\mathbf{0}](1)=0,$
for all $n\in\mathbb{N}$.
There are particular cases in which definitions (LR0), (LR1), (LR2) and (LR3) are all equivalent. We say that a set $\mathcal{D}$ is closed when for all $x\in\mathcal{D},y\notin\mathcal{D}$ and for all $T\in\mathcal{T}$, we have $T(x,y)=0$.
@@
We show that $\text{(LR1)}\Rightarrow\text{(LR3)}$. The thesis will follow from this and Proposition 3.5. We need to prove that, if $x\rightharpoondown y$ under (LR1), then there exists $n\in\mathbb{N}$ such that
@@
for all $T_{1},\dots,T_{n}\in\mathcal{T}$. Let us pick $n=k^{*}$ recalling that, by hypothesis, we have $x\in\mathcal{D}_{k^{*}}$. For all $T_{1}\in\mathcal{T}$, there exists $x_{1}\in\mathcal{D}_{k^{*}-1}$ such that $T_{1}(x,x_{1})>0$. Iterating this argument, we get to $\mathcal{D}$ in at most $k^{*}$ steps. If, for some sequence of matrices, there is a positive probability of hitting $\mathcal{D}$ at time $k^{\prime}<k^{*}$, then, due to the closedness of $\mathcal{D}$, the same would hold at time $k^{*}$. ∎
An alternative proof of the equivalence between (LR1) and (LR3) under the closedness hypothesis can also be found in [4].
@@ 3.2 Lower Hitting Probabilities
Using the definition of lower reachability, we deduce a straightforward relationship between the sets $\mathcal{A}_{\mathcal{T}}$ and $\mathcal{C}_{T}$.
@@
The set $\mathcal{A}_{\mathcal{T}}$ can be rewritten as
$=\{x\in A^{c}\mid\exists T\in\mathcal{T}:x\in\mathcal{C}_{T}\},$
and the claim follows. ∎
We may interpret $\mathcal{A}_{\mathcal{T}}$ as the set of states whose lower probability of reaching $A$ is zero, since there exists a transition matrix that disconnects $\mathcal{A}_{\mathcal{T}}$ from $A$. We are finally able to characterise lower hitting probabilities as the unique solution of a system of equations.
@@
The vector of lower hitting probabilities $\underline{p}^{\mathcal{T}}$ satisfies $\underline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$ and $\underline{p}^{\mathcal{T}}(x)=0$ if and only if $x\in\mathcal{A}_{\mathcal{T}}$. Moreover, $\underline{p}^{\mathcal{T}}$ is the unique solution of
$\underline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\cdot\underline{T}\underline{p}^{\mathcal{T}}.$ (15)
@@
Let us start with the claims about the structure of $\underline{p}^{\mathcal{T}}$. We know from Krak et al. [19, Corollary 19] that
$\underline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\underline{T}\underline{p}^{\mathcal{T}}$
@@
has a minimal nonnegative solution $\underline{p}^{\mathcal{T}}\in[0,1]^{\mathcal{X}}$. By the definition of the system, this solution satisfies $\underline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$. The fact that $\underline{p}^{\mathcal{T}}(x)=0$ if and only if $x\in\mathcal{A}_{\mathcal{T}}$ follows from the definition of lower hitting probability, Lemma 3.10, and Proposition 2.5. Therefore, $\underline{p}^{\mathcal{T}}$ is the minimal nonnegative solution of (15).
Now we prove the uniqueness of $\underline{p}^{\mathcal{T}}$. Let $q\in[0,1]^{\mathcal{X}}$ be a solution of (15). By minimality, we know that $q\geq\underline{p}^{\mathcal{T}}$, therefore we need to show that $q\leq\underline{p}^{\mathcal{T}}$. Since $q$ and $\underline{p}^{\mathcal{T}}$ already agree on $A$ and $\mathcal{A}_{\mathcal{T}}$, it suffices to show that $q|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq\underline{p}^{\mathcal{T}}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$.
By Krak et al. [19, Theorem 18], there exists $T^{*}\in\mathcal{T}$ such that $p^{T^{*}}=\underline{p}^{\mathcal{T}}$. Since $p^{T^{*}}(x)>0$ if and only if $x\in\mathcal{A}_{\mathcal{T}}$, by Proposition 2.5, we have $\mathcal{C}_{T^{*}}=\mathcal{A}_{\mathcal{T}}$ and
$\underline{p}^{\mathcal{T}}=p^{T^{*}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\cdot T^{*}p^{T^{*}}.$
From (15) applied to $q$, we obtain
$q|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=(\underline{T}q)|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq(T^{*}q)|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}.$
Now we apply Lemma 2.6 and we get
$q|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\leq(T^{*}q)|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}=T^{*}|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}q|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}+(T^{*}\mathbb{1}_{A})|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}.$
@@
Recursively expanding this expression for $q|_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}$ yields, after $m\in\mathbb{N}$ expansions
$q|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}\leq(T^{*}|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}})^{m}q|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}+\left[\sum_{k=0}^{m-1}(T^{*}|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}})^{k}\right](T^{*}\mathbb{1}_{A})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}.$
Since $\mathcal{C}_{T^{*}}={\mathcal{A}_{\mathcal{T}}}$, we let $m\to+\infty$ and we apply Lemma 2.12:
$q|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}\leq(I-T^{*}|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}})^{-1}(T^{*}\mathbb{1}_{A})|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=p^{T^{*}}|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}}=\underline{p}^{\mathcal{T}}|_{A^{c}\setminus{\mathcal{A}_{\mathcal{T}}}},$
where the penultimate equality follows from Proposition 2.13. Thus $q=\underline{p}^{\mathcal{T}}$, which is what we wanted to prove. ∎
@@ 3.3 Upper Hitting Probabilities
Similarly to lower hitting probabilities, there is a straightforward connection between sets $\mathscr{W}_{\mathcal{T}}$ and $\mathcal{C}_{T}$.
@@
The set $\mathscr{W}_{\mathcal{T}}$ can be rewritten as
$=\{x\in A^{c}\mid\forall T\in\mathcal{T}:x\in\mathcal{C}_{T}\},$
and the claim follows. ∎
We may interpret $\mathscr{W}_{\mathcal{T}}$ as the set of states whose upper probability of reaching $A$ is zero, since every transition matrix disconnects $\mathscr{W}_{\mathcal{T}}$ to $A$.
The analogue of Proposition 3.11 for upper hitting probabilities fails in general. Nevertheless, we find that $\overline{p}^{\mathcal{T}}$ satisfies $\overline{p}^{\mathcal{T}}(x)=1$ for all $x\in A$ and $\overline{p}^{\mathcal{T}}(x)=0$ if and only if $x\in\mathscr{W}_{\mathcal{T}}$. These properties follow immediately from Krak et al. [19, Corollary 19], the definition of upper hitting probability, and Lemma 3.13. We know from Krak et al. [19, Corollary 19] that $\overline{p}^{\mathcal{T}}$ is the minimal nonnegative solution of $\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\overline{T}\overline{p}^{\mathcal{T}}$, therefore it is also the minimal nonnegative solution of
$\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\cdot\overline{T}\overline{p}^{\mathcal{T}}.$ (17)
Unlike the lower hitting probability, the vector $\overline{p}^{\mathcal{T}}$ need not be the unique solution of (17). Below we present an example in which another vector $q>\overline{p}^{\mathcal{T}}$ also satisfies (17).
Example 3.14.
Let $\mathcal{X}=\{1,2,3\}$ be the state space, and let $\{2\}$ be the target set. Let the set of transition matrices on $\mathcal{X}$ be
$\mathcal{T}=\operatorname{co}\left\{I,T_{n}=\begin{bmatrix}1-\frac{2}{n}&\frac{1}{n}-\frac{1}{n^{2}}&\frac{1}{n}+\frac{1}{n^{2}}\\ 0&1&0\\ 0&0&1\\ \end{bmatrix}:n\geq 2\right\},$
where by “co” we denote the convex hull and by “II” the identity matrix. The situation is represented by the following graph:
[…]
Since $A=\{2\}$ and $\mathscr{W}_{\mathcal{T}}=\{3\}$, we have that, for every $T\in\mathcal{T}$, $p^{T}(2)=1$ and $p^{T}(3)=0$. When starting from state 1{1} and evolving under the transition matrix $T_{n}$, the hitting probability $p^{T_{n}}$ satisfies
$p^{T_{n}}(1)=\left(1-\dfrac{2}{n}\right)p^{T_{n}}(1)+\left(\dfrac{1}{n}-\dfrac{1}{n^{2}}\right).$
Therefore
$p^{T_{n}}(1)=\dfrac{n}{2}\ \dfrac{n-1}{n^{2}}=\dfrac{n-1}{2n}\xrightarrow{n\to\infty}\dfrac{1}{2},$
and $p^{T_{n}}(1)<1/2$ for all $n\geq 2$. Since no transition matrix in $\mathcal{T}$ yields a hitting probability greater or equal than $1/2$, we have that $\overline{p}^{\mathcal{T}}(1)=1/2$. The vector $\overline{p}^{\mathcal{T}}=(1/2,1,0)^{\top}$ satisfies (17):
$1/2=\overline{p}^{\mathcal{T}}(1)=\sup_{T\in\mathcal{T}}\left[T\cdot\begin{pmatrix}1/2\\ 1\\ 0\end{pmatrix}\right](1)=\left[I\cdot\begin{pmatrix}1/2\\ 1\\ 0\end{pmatrix}\right](1)=1/2,$
where the third equality follows from the identity matrix $I\in\mathcal{T}$ attaining the supremum: $\overline{T}\overline{p}^{\mathcal{T}}](1)=[I\overline{p}^{\mathcal{T}}](1)$.
The vector $q=(1,1,0)^{\top}$ satisfies (17) too. This follows from the fact that
$T_{n}q=\begin{bmatrix}1-\frac{2}{n}&\frac{1}{n}-\frac{1}{n^{2}}&\frac{1}{n}+\frac{1}{n^{2}}\\ 0&1&0\\ 0&0&1\\ \end{bmatrix}\begin{pmatrix}1\\ 1\\ 0\end{pmatrix}=\begin{pmatrix}1-\frac{1}{n}-\frac{1}{n^{2}}\\ 1\\ 0\end{pmatrix}<\begin{pmatrix}1\\ 1\\ 0\end{pmatrix}=q,$
and that $\overline{T}q=Iq=q$. More generally, every vector $q_{\alpha}\coloneqq(\alpha,1,0)^{\top}$ with $1/2\leq\alpha\leq 1$ is a solution of (17).
To restore the uniqueness of $\overline{p}^{\mathcal{T}}$ we may impose additional hypotheses; for instance, it suffices that every transition matrix $T\in\mathcal{T}$ has the same set $\mathcal{C}_{T}$, or equivalently $\mathscr{W}_{\mathcal{T}}=\mathcal{A}_{\mathcal{T}}$.
@@
If the set $\mathcal{T}$ is such that $\mathscr{W}_{\mathcal{T}}=\mathcal{A}_{\mathcal{T}}$, then the vector of upper hitting probabilities $\overline{p}^{\mathcal{T}}$ is the unique solution of
$\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\cdot\overline{T}\overline{p}^{\mathcal{T}}.$ (18)